Salah’s net income for the year ended December 31, Year 2 was $197,000. Information from Salah’s comparative balance sheets is given below. Compute the cash paid for dividends during Year 2.

At December 31 Year 2 Year 1
Common Stock, $5 par value $ 512,000 $ 460,800
Paid-in capital in excess of par 960,000 863,800
Retained earnings 700,000 592,800
Answer #1

Cash paid for dividends

= Net income for year 2 - Increase in Retained earnings

= 197000 - (700000-592800=

= 197000 - 107200

= 89800